OpenStudy (anonymous):

The lengths of the sides of a rectangle become 4 times greater than the original measurements. Which statement is true about the rectangle? Its area becomes 4 times greater, and perimeter becomes 20 times greater. Its area becomes 16 times greater, and perimeter becomes 4 times greater. Its area becomes 4 times greater, and perimeter becomes 16 times greater. Its area becomes 20 times greater, and perimeter becomes 4 times greater.

OpenStudy (anonymous):

@Hero @jim_thompson5910 @ash2326

OpenStudy (akashdeepdeb):

Well so let the initial dimensions be L and B So area would be L*B So after they are quadrupled we get 4L and 4B So the area would become 4L*4B = 16LB! [16 TIMES] And new perimeter would become = 2(4L + 4B) = 8(L+B) [4 TIMES] Understood? :)

OpenStudy (anonymous):

Second answer????

OpenStudy (akashdeepdeb):

Spot on! :)

OpenStudy (anonymous):

Thanks

OpenStudy (akashdeepdeb):

:)